1) Yield to maturity: Rudy Sandberg wants to invest in four-year bonds that are currently priced at $868.43. These bonds have a coupon rate of 6 percent and pay semiannual coupon payments. What is the current market yield on this bond?

2) Realized yield: Josh Kavern bought ten-year, 12 percent coupon bonds issued by the U.S. Treasury three years ago at $913.44. If he sells these bonds, which have a face value of $1,000, at the current price of $804.59, what is the realized yield on the bonds? Assume similar coupon-paying bonds make annual coupon payments.

3)Peabody Corp. has seven-year bonds outstanding. The bonds pay a coupon of 8.375 percent semiannually and are currently worth $1,063.49. The bonds can be called in three years at a price of $1,075.

a. What is the yield to maturity on the bond?

b. What is the effective annual yield?

c. What is the yield to call on the bond?

d. If you plan to invest in one of these bonds today, what is the expected yield on the investment? Explain.
Answer #1

1)

Function Arguments ? x RATE 4*2 Nper Pmt 60/2 = 8 = 30 = -868.43 - 1000 1 1 1 -868.43 FV Type 1000 o = 0.050388751 Returns th

Hence, Current market yield is 10.08% (5.04%*2)

*Please rate thumbs up